OpenStudy (anonymous):

How do you find the derivative of 1/(x+7)^-2

OpenStudy (anonymous):

Do chain rule. \[f(x) = \frac{1}{(x+7)^{2}}\] I assume that's the equation you are writing and not \[f(x) = \frac{1}{(x+7)^{-2}}=(x+7)^{2}\] Here's chain rule if you forgot. \[f(x) = a*g(x)^{n}\] \[f'(x) = (n-1)*a*g(x)^{n-1}*g'(x)\]

OpenStudy (anonymous):

Ok so would the answer be 2x+14

OpenStudy (anonymous):

if your \[f(x) = (x+7)^{2}\] then yes

OpenStudy (anonymous):

Thank you!